# Efficient Reinforcement Finetuning via Adaptive Curriculum Learning

### v2

@@ Abstract
Reinforcement finetuning (RFT) has shown great potential for enhancing the mathematical reasoning capabilities of large language models (LLMs), but it is often sample- and compute-inefficient, requiring extensive training. In this work, we introduce AdaRFT (Adaptive Curriculum Reinforcement Finetuning), a method that significantly improves both the efficiency and final accuracy of RFT through adaptive curriculum learning. AdaRFT dynamically adjusts the difficulty of training problems based on the model’s recent reward signals, ensuring that the model consistently trains on tasks that are challenging but solvable. This adaptive sampling strategy accelerates learning by maintaining an optimal difficulty range, avoiding wasted computation on problems that are too easy or too hard. AdaRFT requires only a lightweight extension to standard RFT algorithms like Proximal Policy Optimization (PPO), without modifying the reward function or model architecture. Experiments on competition-level math datasets—including AMC, AIME, and IMO-style problems—demonstrate that AdaRFT significantly improves both training efficiency and reasoning performance. We evaluate AdaRFT across multiple data distributions and model sizes, showing that it reduces the number of training steps by up to 2 and improves accuracy by a considerable margin, offering a more scalable and effective RFT framework.
@@ 2 Related Work
Most RFT pipelines build on Proximal Policy Optimization (PPO) (Schulman et al., 2017b), with recent variants like RAFT (Dong et al., 2023), ReMax (Li et al., 2024), GRPO (DeepSeek-AI et al., 2025), and REINFORCE++ (Hu, 2025), aiming to reduce computational overhead by simplifying RL components. While effective, these methods often trade off stability or sample efficiency. In parallel, data-centric strategies have emerged as promising alternatives for efficient finetuning. LIMO (Ye et al., 2025) and s1 (Muennighoff et al., 2025) show that small, carefully selected supervised datasets can yield strong downstream performance, but their success hinges on manual curation, prompt engineering, and careful dataset construction, which may not generalize across tasks or models. LIMR (Li et al., 2025) and Wang et al. (2025) proposes scoring training examples based on their estimated learning impact, enabling selective finetuning with fewer samples. Yet, computing these scores requires a full training run, and the scores must be recomputed for each new model, limiting practicality and scalability. In contrast, AdaRFT introduces a lightweight, model-agnostic curriculum learning strategy that dynamically adjusts task difficulty based on reward feedback. This allows continuous adaptation to the model’s capabilities, improving convergence speed and final accuracy without modifying the RL algorithm or requiring manual data curation.
@@
Curriculum learning (CL) structures training by presenting tasks in an organized progression, typically from easy to hard, to enhance learning efficiency and generalization (Bengio et al., 2009). In RL, CL methods include task sorting by difficulty (Zaremba & Sutskever, 2015; Justesen et al., 2018; Wang et al., 2019), teacher-student frameworks that adaptively select tasks based on learning progress (Matiisen et al., 2017; Portelas et al., 2019), and self-play approaches that induce automatic curricula through agent competition (Sukhbaatar et al., 2018). Other strategies use intermediate-goal generation in sparse-reward settings (Florensa et al., 2018), unsupervised skill discovery (Jabri et al., 2019), or knowledge transfer via progressive networks and imitation (Czarnecki et al., 2018; Rusu et al., 2022). While CL is well-studied in classical RL, its application to RFT of LLMs is still limited. Existing methods typically use staged training with hand-designed difficulty tiers (Wen et al., 2025; Luo et al., 2025; Song et al., 2025), or online filtering schemes that repeatedly sample and discard data until rewards reach a target range (Bae et al., 2025; Yu et al., 2025). These methods either lack adaptability or introduce significant computational overhead due to repeated rollouts. In contrast, AdaRFT is among the first truly adaptive curriculum learning approaches for RFT: it continuously adjusts task difficulty based on the model’s reward signal, enabling efficient, scalable training without fixed schedules or repeated rollouts.
@@ 3 AdaRFT
Let be a dataset of mathematical problems, each annotated with a precomputed difficulty score . For example, the MATH dataset (Hendrycks et al., 2021) includes preassigned difficulty levels ranging from 1 to 5. These difficulty scores can also be derived through other methods. One common approach is to use human expert annotations; for instance, problems from the International Mathematical Olympiad (IMO) are generally considered more difficult than those from the American Mathematics Competitions (AMC). Alternatively, difficulty can be estimated based on the model’s empirical success rate: problems that are frequently solved correctly receive lower difficulty scores, while problems that are rarely solved correctly are given higher scores. Another approach is to train a separate model, such as a regression or classification model, to estimate difficulty from features of the problem. Given a predefined difficulty for each problem, the objective is to train a policy that improves its problem-solving ability by dynamically adjusting the training curriculum according to the model’s current performance. Our proposed algorithm, AdaRFT, is shown in Algorithm 1.
@@ 3.1 Dynamic Curriculum Sampling
To construct an adaptive curriculum, we define a target difficulty , which represents the current target difficulty level for training. AdaRFT dynamically adjusts based on the model’s reward signal to maintain an optimal difficulty level for learning. At each training step (see Alg. 1), the algorithm computes the absolute difference between the target difficulty and the difficulty of each problem in the dataset (Alg. 1, line 4): for all . The batch of training problems is formed by selecting the problems with the smallest values of (Alg. 1, line 5), producing a batch: . This ensures that the selected problems are closest to the model’s current target difficulty, focusing the learning process on problems that are neither too easy nor too hard.
@@ 3.2 Policy Update
The selected batch is used to train the policy model , which generates responses: . A reward signal is computed based on the correctness of the model’s output (Alg. 1, line 7):
@@ 3.3 Target Difficulty Update
Here, are hyperparameters: is the step size for adjusting the target difficulty, controls the sensitivity of the update, and is the target reward level, representing the desired success rate. The function ensures smooth updates and prevents large jumps in difficulty by saturating for large deviations, while the “clip” function constrains the target difficulty within the valid range . These bounds can be manually specified or automatically derived from the training set, for example by taking the minimum and maximum of the difficulty scores . Intuition and guidance for selecting these hyperparameters are discussed in Section 3.4 and 4.3.
@@ 3.4 Theoretical Justification for Target Reward
where is the model’s success rate on prompt . This implies that the lower bound on the KL divergence, and consequently the gradient magnitude during policy updates, is proportional to the reward variance, which is maximized when . In other words, training on prompts that the model succeeds on roughly half the time may yield the strongest learning signal.
@@ 4.1 Difficulty Estimation
Accurate estimation of problem difficulty is critical for AdaRFT. For difficulty estimation, we use the Qwen 2.5 MATH 7B model (Qwen et al., 2025).111https://huggingface.co/Qwen/Qwen2.5-Math-7B A model that is too strong (e.g., OpenAI o1 (OpenAI et al., 2024b), DeepSeek R1 (DeepSeek-AI et al., 2025)) would solve most problems on the first attempt, leading to poor discrimination between easy and hard problems. Conversely, a model that is too weak (e.g., LLaMA 3.3 1B (Grattafiori et al., 2024)) would fail to solve most problems even after multiple attempts, limiting the signal required for curriculum adaptation.
We select Qwen 2.5 MATH 7B as the estimator because it demonstrates a balanced solving ability. It can solve moderately difficult problems while still failing on more complex problems, thereby providing a more informative difficulty signal. For each problem, the difficulty score is computed as:
where is the number of attempts per problem. In our setup, we use . This approach ensures that difficulty scores reflect the true problem-solving capability of a moderately capable model, preventing over- or under-estimation of difficulty.
@@
To further validate the reliability of our difficulty estimates, we examined their alignment with the difficulty levels provided in the MATH dataset. The MATH dataset comprises 12,500 competition-level mathematics problems sourced from contests such as AMC 10, AMC 12, and AIME. Each problem is categorized into one of five difficulty levels, following the classification system used by the Art of Problem Solving (AoPS) community.222https://artofproblemsolving.com/wiki/index.php/AoPS_Wiki%3ACompetition_ratings In this system, level 1 denotes the easiest problems, while level 5 represents the most difficult. For example, the initial problems in an AMC 8 exam are typically labeled as level 1, whereas AIME problems are often labeled as level 5. As shown in Figure 1(b), there is a clear downward trend in the average solve rate as the labeled difficulty level increases, ranging from 86.0% at level 1 to 52.7% at level 5. Specifically, the AoPS-derived difficulty levels yield a Pearson correlation of () with model success rates. This negative correlation indicates that the model’s empirical performance aligns well with the intended difficulty stratification, reinforcing the utility of both the labeled difficulty levels and our estimation approach in guiding curriculum learning. To further streamline the difficulty estimation process, we also prompted GPT-4o (gpt-4o-0806) (OpenAI et al., 2024a) to assign difficulty levels to the DeepScaleR dataset based on the AoPS rubric. Each problem was presented to GPT-4o with a request to rate its difficulty according to AoPS guidelines (the full prompt is shown in Appendix A.3). This approach provides a lightweight and scalable alternative to rollout-based estimation. As shown in Figure 1(b), GPT-4o’s difficulty ratings also correlate well with the model success rates, with a Pearson correlation of (), making it a practical proxy for curriculum scheduling when computational resources are constrained.
@@ 4.2 Dataset
We use the DeepScaleR dataset (Luo et al., 2025) as the training set. DeepScaleR compiles problems from multiple sources, including the American Invitational Mathematics Examination (AIME) from 1984 to 2023 and the American Mathematics Competitions (AMC) prior to 2023. The dataset also includes problems from the Omni-MATH (Gao et al., 2024) and Still datasets (Team, 2025), which feature problems from various national and international math competitions. This results in a diverse and challenging training set, covering a wide range of mathematical domains and difficulty levels.
@@ 4.3 Training Setup
We trained two models on the three difficulty-based distributions of the DeepScaleR dataset described in Section 4.2: Qwen 2.5 7B333https://huggingface.co/Qwen/Qwen2.5-7B and Qwen 2.5 MATH 1.5B444https://huggingface.co/Qwen/Qwen2.5-Math-1.5B. This setup allows us to evaluate the effectiveness of AdaRFT on models with different initial performance levels when exposed to skew-hard, skew-easy, and uniform problem distributions.
All models were trained using three different approaches: (1) the standard PPO algorithm, (2) AdaRFT (PPO), our method that integrates adaptive curriculum learning with PPO (see Section 3), and (3) PPO with filtered data, a strong baseline that trains PPO using filtered data based on pass@ accuracy. Following the approach of prior work (Bae et al., 2025; Hu et al., 2025; Zyphra, 2025), we perform a pass@40 analysis for each model and data distribution combination. For the data filtering baseline (3), we remove examples that are either too easy or too hard, excluding problems with solved percentages or . This focuses training on problems of intermediate difficulty. However, this filtering discards a significant portion of the training data, including many potentially useful examples. Moreover, because difficulty is determined via pass@ metrics, filtering must be recomputed each time the model or data distribution changes. In contrast, AdaRFT only requires difficulty estimation once per dataset and adapts to any model during training. For AdaRFT, the target difficulty was dynamically adjusted throughout training based on the model’s reward signal. This adjustment ensured that the model consistently encountered problems that matched its current skill level, preventing the model from being overwhelmed by overly difficult problems or stagnating on problems that were too easy.
@@
The training batch size was set to , with the target reward set to 0.5 to promote learning at a balanced success rate. The sensitivity parameter and step size were tuned using a validation set to ensure stable curriculum updates. We set , , and the initial target difficulty . The step size acts as a scaling factor between the reward signal and the difficulty metric. Since the difficulty metric ranges from 0 to 100 and the reward ranges from 0 to 1, a target reward implies that the maximum reasonable adjustment to the difficulty metric should be around 50. Therefore, we set to scale the reward signal appropriately to the difficulty range. The sensitivity parameter controls the slope of the function. Setting to 2 makes the function behave approximately linearly when the difference between the average reward and the target reward is small. The intuition behind using the function is that when the average reward is close to the target reward, a roughly linear adjustment is appropriate. However, when the average reward deviates significantly from the target reward, linear adjustments may be too large, leading to instability. The function smooths out these adjustments, allowing for more controlled changes when the difference is large while maintaining sensitivity when the difference is small. Both models were trained on 8 A100 GPUs for approximately 100 steps. The implementation details can be found in Appendix A.
5 Result

### v3

@@ Abstract
Reinforcement finetuning (RFT) has shown great potential for enhancing the mathematical reasoning capabilities of large language models (LLMs), but it is often sample- and compute-inefficient, requiring extensive training. In this work, we introduce AdaRFT (Adaptive Curriculum Reinforcement Finetuning), a method that significantly improves both the efficiency and final accuracy of RFT through adaptive curriculum learning. AdaRFT dynamically adjusts the difficulty of training problems based on the model’s recent reward signals, ensuring that the model consistently trains on tasks that are challenging but solvable. This adaptive sampling strategy accelerates learning by maintaining an optimal difficulty range, avoiding wasted computation on problems that are too easy or too hard. AdaRFT requires only a lightweight extension to standard RFT algorithms like Proximal Policy Optimization (PPO), without modifying the reward function or model architecture. Experiments on competition-level math datasets demonstrate that AdaRFT significantly improves both training efficiency and reasoning performance. We evaluate AdaRFT across multiple data distributions and model sizes, showing that it reduces training time by up to 2 and improves accuracy by a considerable margin, offering a more scalable and effective RFT framework.
Efficient Reinforcement Finetuning via Adaptive Curriculum Learning
Taiwei Shi†, Yiyang Wu†, Linxin Song†, Tianyi Zhou▽, Jieyu Zhao†
†University of Southern California, ▽University of Maryland, College Park
{taiweish, wuwendy, linxinso, jieyuz}@usc.edu, tianyi@umd.edu
Code: github.com/limenlp/verl
Dataset: huggingface.co/datasets/lime-nlp/DeepScaleR_Difficulty
@@ 2 Related Work
Most RFT pipelines build on Proximal Policy Optimization (PPO) (Schulman et al., 2017b), with recent variants like RAFT (Dong et al., 2023), ReMax (Li et al., 2024), GRPO (DeepSeek-AI et al., 2025), and REINFORCE++ (Hu, 2025), aiming to reduce computational overhead by simplifying RL components. While effective, these methods often trade off stability or sample efficiency. In parallel, data-centric strategies have emerged as promising alternatives for efficient finetuning. LIMO (Ye et al., 2025) and s1 (Muennighoff et al., 2025) show that small, carefully selected supervised datasets can yield strong downstream performance, but their success hinges on manual curation, prompt engineering, and careful dataset construction, which may not generalize across tasks or models. LIMR (Li et al., 2025) and Wang et al. (2025) proposes scoring training examples based on their estimated learning impact, enabling selective finetuning with fewer samples. Yet, computing these scores requires a full training run, and the scores must be recomputed for each new model, limiting practicality and scalability. Moreover, reducing the number of training samples does not inherently translate to improved efficiency. Models still require a comparable number of optimization steps and wall-clock time to converge. In contrast, AdaRFT introduces a lightweight, model-agnostic curriculum learning strategy that dynamically adjusts task difficulty based on reward feedback. This allows continuous adaptation to the model’s capabilities, improving convergence speed and final accuracy without modifying the RL algorithm or requiring manual data curation.
@@
Curriculum learning (CL) structures training by presenting tasks in an organized progression, typically from easy to hard, to enhance learning efficiency and generalization (Bengio et al., 2009). In RL, CL methods include task sorting by difficulty (Zaremba and Sutskever, 2015; Justesen et al., 2018; Wang et al., 2019), teacher-student frameworks that adaptively select tasks based on learning progress (Matiisen et al., 2017; Portelas et al., 2019), and self-play approaches that induce automatic curricula through agent competition (Sukhbaatar et al., 2018; Zhao et al., 2025). Other strategies use intermediate-goal generation in sparse-reward settings (Florensa et al., 2018), unsupervised skill discovery (Jabri et al., 2019), or knowledge transfer via progressive networks and imitation (Czarnecki et al., 2018; Rusu et al., 2022). While CL is well-studied in classical RL, its application to RFT of LLMs is still limited. Existing methods typically use staged training with hand-designed difficulty tiers (Wen et al., 2025; Luo et al., 2025; Song et al., 2025), or online filtering schemes that repeatedly sample and discard data until rewards reach a target range (Bae et al., 2025; Yu et al., 2025). These methods either lack adaptability or introduce significant computational overhead due to repeated rollouts. In contrast, AdaRFT is among the first truly adaptive curriculum learning approaches for RFT: it continuously adjusts task difficulty based on the model’s reward signal, enabling efficient, scalable training without fixed schedules or repeated rollouts.
@@ 3 AdaRFT
Let be a dataset of mathematical problems, each annotated with a precomputed difficulty score . The score can be either human-annotated or model-estimated. The objective is to train a policy that improves its problem-solving ability by dynamically adjusting the training curriculum according to the model’s current performance. Our proposed algorithm, AdaRFT, is shown in Algorithm 1.
@@ 3.1 Dynamic Curriculum Sampling
To construct an adaptive curriculum, we define a target difficulty , which represents the current target difficulty level for training (more in Section 3.3). AdaRFT dynamically adjusts based on the model’s reward signal to maintain an optimal difficulty level for learning. At each step, the algorithm computes the absolute difference between the target difficulty and the difficulty of each problem in the dataset (Alg. 1, line 4): for all . The batch of training problems is formed by selecting the problems with the smallest values of (Alg. 1, line 5), producing a batch: . This ensures that the selected problems are closest to the model’s current target difficulty, focusing the learning process on problems that are neither too easy nor too hard.
@@ 3.2 Policy Update
The selected batch is used to train the policy model , which generates responses: . A reward signal is computed based on the correctness of the model’s output (Alg. 1, line 7): if the response is correct, and if the response is incorrect. The average reward over the batch is computed as (Alg. 1, line 7): . The policy can then be updated using a reinforcement learning algorithm such as PPO, GRPO, or REINFORCE++ (Alg. 1, line 8): .
@@ 3.3 Target Difficulty Update
The full derivation is provided in Appendix B.2. This update rule is a smooth and stabilized variant of a standard linear mapping between reward space and difficulty space. Here, are hyperparameters: is the step size for adjusting the target difficulty, controls the sensitivity of the update, and is the target reward level, representing the desired success rate. The function ensures smooth updates and prevents large jumps in difficulty by saturating for large deviations, while the “clip” function constrains the target difficulty within the valid range . These bounds can be manually specified or automatically derived from the training set, for example, by taking the minimum and maximum of the difficulty scores . Intuition and guidance for selecting these hyperparameters are discussed in Section 3.4, 4.3, and Appendix B.2.
@@ 3.4 Theoretical Justification for Target Reward
where is the model’s success rate on prompt . This implies that the lower bound on the KL divergence, and consequently the gradient magnitude during policy updates, is proportional to the reward variance, which is maximized when . In other words, training on prompts that the model succeeds on roughly half the time may yield the strongest learning signal. In Section 5 and Appendix A.2, we conduct an ablation study by varying the target reward , demonstrating that setting consistently leads to the best performance, supporting the hypothesis that training on prompts with a success rate near 50% provides the most informative learning signal.
@@ 4.1 Difficulty Estimation
Accurate estimation of problem difficulty is critical for AdaRFT. For difficulty estimation, we select the Qwen 2.5 MATH 7B model (Qwen et al., 2025) because it demonstrates a balanced solving ability. A model that is too strong (e.g., OpenAI o1 (OpenAI et al., 2024b), DeepSeek R1 (DeepSeek-AI et al., 2025)) would solve most problems on the first attempt, leading to poor discrimination between easy and hard problems. Conversely, a model that is too weak (e.g., LLaMA 3.3 1B (Grattafiori et al., 2024)) would fail to solve most problems even after multiple attempts, limiting the signal required for curriculum adaptation. For each problem, the difficulty score is computed as: , where denotes the number of successful attempts on problem , and is the number of attempts per problem. In our setup, we use .
@@
To further validate the reliability of our difficulty estimates, we examined their alignment with the difficulty levels provided in the MATH dataset. The MATH dataset comprises 12,500 competition-level mathematics problems sourced from contests such as the American Mathematics Competitions (AMC) and the American Invitational Mathematics Examination (AIME). Each problem is categorized into one of five difficulty levels, following the classification system used by the Art of Problem Solving (AoPS) community (details in Appendix B.3). In this system, level 1 denotes the easiest problems, while level 5 represents the most difficult. As shown in Figure 1(b), there is a clear downward trend in the average solve rate as the labeled difficulty level increases, ranging from 86.0% at level 1 to 52.7% at level 5. Specifically, the AoPS-derived difficulty levels yield a Pearson correlation of () with model success rates. This negative correlation indicates that the model’s empirical performance aligns well with the intended difficulty stratification, reinforcing the utility of both the labeled difficulty levels and our estimation approach in guiding curriculum learning. To further streamline the difficulty estimation process, we also prompted GPT-4o (gpt-4o-0806) (OpenAI et al., 2024a) to assign difficulty levels to the DeepScaleR dataset based on the AoPS rubric. Each problem was presented to GPT-4o with a request to rate its difficulty according to AoPS guidelines (the full prompt is shown in Appendix B.3). This approach provides a lightweight and scalable alternative to rollout-based estimation. As shown in Figure 1(b), GPT-4o’s difficulty ratings also correlate well with the model success rates, with a Pearson correlation of (), making it a practical proxy for curriculum scheduling when computational resources are constrained.
@@ 4.2 Dataset
We use the DeepScaleR dataset (Luo et al., 2025) as the training set. DeepScaleR compiles problems from multiple sources, including AIME from 1984 to 2023 and AMC prior to 2023. The dataset also includes problems from the Omni-MATH (Gao et al., 2024) and Still datasets (Team, 2025), which feature problems from various national and international math competitions. This results in a diverse and challenging training set, covering a wide range of mathematical domains and difficulty levels.
@@ 4.3 Training Setup
We trained two models on the three difficulty-based distributions of the DeepScaleR dataset described in Section 4.2: Qwen 2.5 7B and Qwen 2.5 MATH 1.5B. This setup allows us to evaluate the effectiveness of AdaRFT on models with different initial performance levels when exposed to skew-difficult, skew-easy, and uniform problem distributions. All models were trained using four different approaches: (1) the standard PPO algorithm, (2) AdaRFT (PPO), our method that integrates adaptive curriculum learning with PPO (see Section 3), (3) PPO with filtered data, a baseline that trains PPO on data filtered by pass@ accuracy, and (4) PPO with a fixed curriculum schedule.
For the data filtering baseline (3), following prior work (Bae et al., 2025; Hu et al., 2025; Zyphra, 2025), we first run a pass@40 analysis for each combination of model and data distribution. We then discard examples that are either too easy or too hard, removing all problems with solved rates or . This restricts training to problems of intermediate difficulty. However, this procedure removes a large fraction of the data, including many potentially informative examples. In addition, because difficulty is defined using pass@ metrics, the filtering must be recomputed whenever the model or the data distribution changes.
For the fixed curriculum baseline (4), we follow the approach of prior work (Parashar et al., 2025; Team et al., 2025). In this setting, the difficulty of sampled problems follows a predetermined schedule that increases linearly over training steps. Suppose training runs for total steps. We then define a target difficulty at step by
and at each step we sample problems whose estimated difficulty matches this target. Unlike AdaRFT, this schedule increases difficulty at a fixed rate regardless of how quickly or slowly the model learns.
@@
The training batch size was set to , with the target reward set to 0.5 to promote learning at a balanced success rate. The sensitivity parameter and step size were tuned using a validation set to ensure stable curriculum updates. We set , , and the initial target difficulty . The step size acts as a scaling factor between the reward signal and the difficulty metric. Since the difficulty metric ranges from 0 to 100 and the reward ranges from 0 to 1, a target reward implies that the maximum reasonable adjustment to the difficulty metric should be around 50. Therefore, we set to scale the reward signal appropriately to the difficulty range. The sensitivity parameter controls the slope of the function. Setting to 2 makes the function behave approximately linearly when the difference between the average reward and the target reward is small. The function smooths out the adjustments, allowing for more controlled changes when the difference is large while maintaining sensitivity when the difference is small. Both models were trained on 8 A100 GPUs for approximately 100 steps. The implementation details can be found in Appendix B.
5 Results and Analysis
@@
| Model | Setup | Method | Final Acc (%) | Avg Step Time (s) | Extra Steps (%) | Extra Steps | Extra Time (s) |
| Qwen2.5 Math 1.5B | skew-difficult | AdaRFT | 40.48 | 122.24 | 0.0% | +0 | 0.00 |
| PPO | 37.41 | 132.95 | 71.7% | +43 | 5716.85 | ||
| PPO (w/ Filter) | 37.22 | 128.20 | 81.7% | +49 | 6281.80 | ||
| PPO (w/ FC) | 37.22 | 130.91 | 26.7% | +16 | 2094.56 | ||
| uniform | AdaRFT | 41.11 | 121.31 | 0.0% | +0 | 0.00 | |
| PPO | 37.20 | 126.82 | 56.7% | +34 | 4311.88 | ||
| PPO (w/ Filter) | 37.87 | 126.35 | 86.7% | +52 | 6570.20 | ||
| PPO (w/ FC) | 37.27 | 126.40 | 80.0% | +48 | 6067.20 | ||
| skew-easy | AdaRFT | 39.18 | 120.52 | 0.0% | +0 | 0.00 | |
| PPO | 38.46 | 121.15 | 26.7% | +16 | 1938.40 | ||
| PPO (w/ Filter) | 38.15 | 115.12 | 35.0% | +21 | 2417.52 | ||
| PPO (w/ FC) | 38.09 | 121.99 | 58.3% | +35 | 4269.65 | ||
| Qwen2.5 7B | skew-difficult | AdaRFT | 46.83 | 239.92 | 0.0% | +0 | 0.00 |
| PPO | 44.17 | 246.21 | 60.0% | +24 | 5909.04 | ||
| PPO (w/ Filter) | 45.35 | 254.22 | 62.5% | +25 | 6355.50 | ||
| PPO (w/ FC) | 45.54 | 243.12 | 22.5% | +9 | 2188.08 | ||
| uniform | AdaRFT | 46.92 | 234.16 | 0.0% | +0 | 0.00 | |
| PPO | 44.70 | 243.82 | 32.5% | +13 | 3169.66 | ||
| PPO (w/ Filter) | 44.57 | 263.11 | 57.5% | +23 | 6051.53 | ||
| PPO (w/ FC) | 44.32 | 240.62 | 17.5% | +7 | 1684.34 | ||
| skew-easy | AdaRFT | 45.94 | 247.44 | 0.0% | +0 | 0.00 | |
| PPO | 45.07 | 235.27 | 50.0% | +20 | 4705.40 | ||
| PPO (w/ Filter) | 44.98 | 233.13 | 42.5% | +17 | 3963.21 | ||
| PPO (w/ FC) | 43.46 | 240.66 | 57.5% | +23 | 5535.18 |